### v1

# Measurement-Induced Spectral Transition

Abstract
We show that noisy quantum dynamics exposed to weak measurements exhibit a spectral transition between gapless and gapped phases. To this end, we employ the Lyapunov spectrum obtained through singular values of a non-unitary matrix describing the dynamics. We discover that the gapless and gapped phases respectively correspond to the volume-law and area-law phases of the entanglement entropy for the dominant Lyapunov vector. This correspondence between the spectral gap and the scaling of entanglement offers an intriguing analogy with ground-state phase transitions. We also discuss some crucial differences from ground-state transitions, such as the scaling law of the entanglement and the exponentially small gaps. Furthermore, we show that the spectral transition leads to the transition of the timescale for the memory loss of initial states.
Introduction.–
In a quantum system described by a time-independent generator, such as the Hamiltonian for isolated systems and the Liouvillian for dissipative systems, the generator’s spectrum and gap provide essential information about the system. For example, quantum phase transitions of ground states are accompanied by the gap closing of the Hamiltonian [1, 2], and the Liouvillian gap is related to the asymptotic decay rate towards the stationary state [3, 4, 5, 6, 7, 8, 9]. For the quantum phase transitions, ground states in gapless regimes exhibit behaviors qualitatively different from those in gapped phases in terms of the entanglement entropy and correlation functions [10, 11, 12, 13].
The entanglement entropy, which captures intrinsic features of quantum states, has also attracted recent huge attention in noisy dynamics of monitored quantum systems, where generators depend on time randomly. Owing to the competition between unitary dynamics and quantum measurements, where the former and latter respectively enhance and suppress the entanglement growth, measurement-induced entanglement transitions occur [14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42]. Many intriguing aspects of such measurement-induced transitions have been explored, including critical properties [14, 17, 18, 21, 23, 25, 34, 41, 42], timescales of relaxation [22, 28, 29, 38, 43], and multifractality [33].
Then, a natural question is whether the measurement-induced phase transitions under noisy dynamics can be understood as a spectral transition as for the ground-state phase transition in equilibrium. Unfortunately, this problem has been elusive due to the intrinsic randomness in the quantum measurement process.
Figure 1: (a) Schematic figure of the spectral transition. With varying a parameter $\eta$, the transition occurs between the gapless and gapped phases, where $\{\varepsilon_{i}(\eta)\}$ are the Lyapunov spectra of the time-evolution matrix. The transition point of the spectral transition coincides with that of the entanglement transition between the volume-law and area-law phases for the dominant Lyapunov vector corresponding to $\varepsilon_{1}(\eta)$. (b) Our model for the dynamics of qubits under measurements. After local unitary gates are applied, quantum measurements with errors $1-\eta$ are carried out for all sites.
In this Letter, we discover a measurement-induced spectral transition leading to the entanglement transition, in analogy with the ground-state phase transitions between gapless and gapped phases. While our monitored system lacks a time-independent generator, we here focus on the Lyapunov spectrum obtained through singular values of a random non-unitary matrix describing the dynamics. As a result, we find a spectral transition from the gapless phase to the gapped phase, as shown in Fig. 1 (a). The gapless and gapped phases respectively correspond to the volume-law and area-law phases of the entanglement entropy for the dominant Lyapunov vector after long times. Such correspondence between the entanglement scaling and the spectral gap is similar to the entanglement transitions of ground states [10, 11, 12, 13]. Meanwhile, in gapless phases, we find that the volume-law of entanglement entropy and exponentially small spectral gaps of monitored dynamics are distinct from the logirithmic-law of entanglement entropy and polynomially small spectral gaps of ground states. We further argue that the spectral transition leads to the transition of the timescale at which observables become independent of initial states, reminiscent of the purification transition.
Setup.–
We consider quantum dynamics depicted in Fig. 1 (b), where $L$ qubits are arrayed on a one-dimensional chain. The local unitary gates, which generate entanglement between nearest neighbor qubits, are given by $4\times 4$ Haar-random unitary matrices. These gates are arranged in a brick-wall manner under the open boundary condition, where the unitary matrix at each time $t$ is denoted as $U_{t}$. The weak quantum measurements at time $t$, described by Kraus operators $M_{t}(\eta)=\bigotimes_{\ell=1}^{L}\mathsf{M}_{\omega_{t,\ell}}(\eta)$ with
$\mathsf{M}_{\pm}(\eta)=\frac{\sigma_{0}\pm\eta\sigma_{3}}{\sqrt{2(1+\eta^{2})}},$ (1)
include the error $1-\eta$ where $0\leq\eta\leq 1$. Here, $\sigma_{0}$ is the $2\times 2$ identity matrix, $\sigma_{3}$ is the $z$ component of the Pauli matrices, $\ell$ represents a position of a qubit, and $\omega_{t,\ell}$ takes $+$ or --, which correspond to the measurement outcomes for $\sigma_{3}^{\ell}$ under the error with $\sigma_{3}^{\ell}$ being the Pauli matrix at $\ell$. The probability distribution of $\{\omega_{t,\ell}\}$ is determined through the Born rule. The parameter $\eta$ describes the strength of the measurement, where $\eta=1$ and $\eta=0$ correspond to the projective measurement and no measurement, respectively. The dynamics from an initial pure state |$\ket{\psi_{0}}$ becomes |$\ket{\psi_{t}(\eta)}=V_{t}(\eta)\ket{\psi_{0}}/\sqrt{\bra{\psi_{0}}V^{\dagger}_{t}(\eta)V_{t}(\eta)\ket{\psi_{0}}}$, where $V_{t}(\eta)=M_{t}(\eta)U_{t}M_{t-1}(\eta)U_{t-1}\cdots M_{1}(\eta)U_{1}$ implicitly depends on measurement outcomes $\{\omega_{t,\ell}\}$.
We define an effective “Hamiltonian,”
$K_{t}(\eta)=-\frac{1}{2t}\ln\left[V_{t}(\eta)V_{t}^{\dagger}(\eta)\right],$ (2)
which leads to an analogy between ground-state phase transitions and measurement-induced transitions. The eigenvalues of $K_{t}(\eta)$, $\{\varepsilon_{t,i}(\eta,L)\}$, are arrayed as $\varepsilon_{t,i}(\eta,L)\leq\varepsilon_{t,i+1}(\eta,L)$ with $1\leq i\leq N:=2^{L}$, where the singular values of $V_{t}(\eta)$ become $\Lambda_{t,i}(\eta)=e^{-\varepsilon_{t,i}(\eta)t}$. The $i$th Lyapunov vector |$\ket{\Psi_{t,i}(\eta)}$ is the eigenmode of $K_{t}(\eta)$ whose eigenvalue is $\varepsilon_{t,i}(\eta)$. The dominant Lyapunov vector |$\ket{\Psi_{t,1}(\eta)}$ corresponds to the ground state of the effective Hamiltonian.
As an indicator for the measurement-induced phase transition, we focus on the spectral gap,
$\Delta(\eta,L)=\varepsilon_{2}(\eta,L)-\varepsilon_{1}(\eta,L),$ (3)
where $\varepsilon_{i}(\eta)=\lim_{t\rightarrow\infty}\varepsilon_{t,i}(\eta)$. When $t$ is sufficiently large such that $\Delta_{t}(\eta,L)=\varepsilon_{t,2}(\eta,L)-\varepsilon_{t,1}(\eta,L)\simeq\Delta(\eta,L)$, the spectral gap $\Delta(\eta,L)$ gives the relaxation time
$\tau_{\delta}(\eta,L)=\left|\frac{\ln(\delta)}{\Delta(\eta,L)}\right|,$ (4)
at which |$\ket{\Psi_{t,i}(\eta)}$ with $i\geq 2$ become negligible within the precision $\delta$ and thus |$\ket{\psi_{t}(\eta)}\simeq\ket{\Psi_{t,1}(\eta)}$ is realized. The dominant Lyapunov vector is similar to the stationary states in open quantum systems described by static generators, in the sense that they are realized in the long-time regime irrespective of initial states. On the other hand, they exhibit distinct behaviors; |$\ket{\Psi_{t,1}(\eta)}$ randomly fluctuates in time, while stationary states are independent of time.
We note that, in several settings, the spectral gap has been computed at the transition point to extract critical properties of measurement-induced transitions [34, 41, 42]. In contrast, we explore the scaling of $\Delta(\eta,L)$ and eigenstate properties in a wide parameter region of $\eta$.
Figure 2: (a) The spectral gap $\Delta(\eta,L)$ as functions of $\eta$ for various system sizes $L$. $\Delta(\eta,L)$ rapidly decreases with increasing $L$ for small $\eta$, while it is almost constant for large $\eta$. (b) The spectral gap $\Delta(\eta)$ in the thermodynamic limit. The spectral transition between the gapless and gapped phases occurs around $\eta_{\mathrm{c}}=0.2$. (c) The entanglement entropy $S^{L/2}(\eta)$ of the dominant Lyapunov vector |$\ket{\Psi_{t,1}(\eta)}$ for various error parameters $\eta$ and system sizes $L$. The inset shows $S^{L/2}(\eta)$ as functions of the system size $L$, where the gapless and gapped phases exhibit the volume-law (orange squares) and the area-law (blue circles) scalings of $S^{L/2}(\eta)$. (d) The mutual information averaged over $T=10^{5}$ steps after $\tau\geq\tau_{\delta}(\eta,L)$. There is a peak around $\eta_{\mathrm{c}}$, corresponding to the entanglement transition between the volume-law and area-law phases. The thresholds of the spectral transition and the entanglement transition coincide.
ground-state transition in equilibrium measurement-induced transition in noisy dynamics
gapped phase gap: $O(L^{0})$,  entanglement entropy: $O(L^{0})$ gap: $O(L^{0})$,  entanglement entropy: $O(L^{0})$
gapless phase gap: $O[1/\mathrm{poly}(L)]$,  entanglement entropy: $O[\ln(L)]$ gap: $O(e^{-L})$,  entanglement entropy: $O(L)$
Table 1: Comparison between ground-state transitions in one-dimensional isolated quantum systems and the measurement-induced transition explored in this work, in terms of the spectral gap and the half-chain entanglement entropy. Behaviors in gapped phases are qualitatively similar, while those in gapless phases are distinct.
Transition of the spectral gap.–
Our monitored dynamics exhibit a transition in terms of the Lyapunov spectrum. Figure 2 (a) shows spectral gaps $\Delta(\eta,L)$ for various $\eta$ and $L$. The details of the numerical calculations are given in Supplemental Material [44]. We can understand that $\Delta(\eta,L)$ rapidly decreases with increasing $L$ for small $\eta$, which indicates $\Delta(\eta)=0$, where
$\Delta(\eta)=\lim_{L\rightarrow\infty}\Delta(\eta,L).$ (5)
In contrast, $\Delta(\eta,L)$ is almost independent of $L$ for large $\eta$, which indicates $\Delta(\eta)\neq 0$. To compute $\Delta(\eta)$, we extrapolate $\Delta(\eta,L)$ using the fitting function $\Delta_{\mathrm{fit}}(\eta,L)=\Delta(\eta)+\alpha(\eta)[\beta(\eta)]^{-L}$. Through the least-square fitting with $10\leq L\leq 22$, we obtain Fig. 2 (b), which clearly shows the transition between the gapless phase with $\Delta(\eta)=0$ and the gapped phase with $\Delta(\eta)\neq 0$. Indeed, for $\eta<\eta_{\mathrm{c}}=0.2$ ($\eta\geq\eta_{\mathrm{c}}$), we see that the bottom value of the error bar is smaller (larger) than 00, indicating that the system is in the gapless (gapped) phase. The spectral transition originates from the competition between unitary dynamics and quantum measurements since the transition vanishes when there is no unitary gate [44]. We note that Refs. [45, 46] explore Lyapunov exponents and purification times of dynamics evolved by non-local matrices, where no transition is found. This indicates that the locality of $U_{t}$ and $M_{t}(\eta)$ plays a key role in measurement-induced transitions, which is also pointed out in Refs. [30, 31].
Ground states in isolated quantum systems also exhibit the gapless-gapped transition, and the behaviors of the gaps in isolated systems and our monitored dynamics are qualitatively similar in gapped phases, as written in Table 1. Meanwhile, there is a qualitative difference in gapless phases; our monitored dynamics exhibit $\Delta(\eta,L)\propto\exp[-L/\xi(\eta)]$ with $\xi(\eta)=1/\ln[\beta(\eta)]$, in contrast to spectral gaps of local Hamiltonians in the critical phase obeying $1/\mathrm{poly}(L)$ [47].
Transition of the entanglement.–
We show that the above spectral transition corresponds to the entanglement transition. We present the entanglement transition for the dominant Lyapunov vector |$\ket{\Psi_{t,1}(\eta)}$ to make a clear comparison with ground-state phase transitions. Note that the correspondence also applies to the entanglement transition for states after $t\propto L$, often discussed in previous works [18], while the data is not shown.
We show that the threshold around $\eta_{\mathrm{c}}=0.2$ of the spectral transition corresponds to that of the entanglement entropy for |$\ket{\Psi_{t,1}(\eta)}$. To this end, we compute the entanglement entropy averaged over $T=10^{4}$ time steps after $\tau\geq\tau_{\delta}(\eta,L)$ with $\delta=10^{-2}$, $S^{A}(\eta)=\frac{1}{T}\sum_{t=\tau+1}^{\tau+T}S_{t}^{A}(\eta)$, where $S_{t}^{A}(\eta)=-\mathrm{tr}_{A}\left(\rho_{t}^{A}(\eta)\ln[\rho_{t}^{A}(\eta)]\right)$ with $\rho_{t}^{A}(\eta)=\mathrm{tr}_{\overline{A}}\left(\ket{\psi_{t}(\eta)}\bra{\psi_{t}(\eta)}\right)$. Here, $\mathrm{tr}_{\bar{A}}$ is the partial trace with respect to the complement of the subsystem $A$. Figure 2 (c) shows the entanglement entropy as functions of $\eta$ for various $L$. We can see that the entanglement entropy of the half chain exhibits the volume-law scaling, $S^{L/2}(\eta)\propto L$, for small $\eta$. On the other hand, the entanglement entropy exhibits the area-law scaling for large $\eta$, $S^{L/2}(\eta)\propto L^{0}$.
It is known that the transition point between the volume-law and area-law phases exhibits the peak of the mutual information $I_{t}^{A,B}(\eta)=S_{t}^{A}(\eta)+S_{t}^{B}(\eta)-S_{t}^{AB}(\eta)$ for |$\ket{\psi_{t}(\eta)}$ [17]. Figure 2 (d) shows the mutual information of |$\ket{\psi_{t}(\eta)}\simeq\ket{\Psi_{t,1}(\eta)}$ averaged over $T=10^{5}$ time steps, $I^{A,B}(\eta)=\frac{1}{T}\sum_{t=\tau+1}^{\tau+T}I_{t}^{A,B}(\eta)$, where $A$ and $B$ are qubits at $\ell=1$ and $\ell=L$, respectively. We can see that $I^{1,L}(\eta)$ exhibits a peak around $\eta_{\mathrm{c}}=0.2$, which coincides with the spectral transition. We note that a peak of the mutual information is also reported in Ref. [17] while the boundary condition differs from that in the present work.
The coincidence between thresholds of the spectral transition and the entanglement transition is analogous to ground-state phase transitions between gapless and gapped phases in equilibrium [10, 11, 12, 13]. In particular, behaviors of the entanglement entropy in gapped phases are qualitatively similar, as listed in Table 1. This suggests that the relation between the spectral gap and the scaling of the entanglement also applies to non-equilibrium phenomena, as well as the equilibrium physics. Meanwhile, |$\ket{\Psi_{t,1}(\eta)}$ of the gapless monitored dynamics with $\eta<\eta_{\mathrm{c}}$ and the ground states of gapless Hamiltonians have a crucial difference; the entanglement exhibits the volume-law scaling in the former case and the logarithmic scaling in the latter case [48]. We note that gapless phases in equilibrium exhibit criticality, while critical behaviors of measurement-induced transitions are observed at the transition point [14, 17, 18, 21, 23, 25, 34, 41, 42], which may be the key to understand the distinction in gapless phases.
Figure 3: Expectation values of $X=\sum_{\ell}\sigma_{1}^{\ell}$ with (a) $\eta=0.1$ in the gapless phase and (b) $\eta=0.3$ in the gapped phase. The black broken lines represent $\tau_{\delta}(\eta,L)=|\ln(\delta)/\Delta(\eta,L)|$ with $\delta=10^{-2}$. Here, different symbols correspond to randomly chosen different initial states, and the same $V_{t}(\eta)$ is applied to all initial states. The expectation values become almost the same after $\tau_{\delta}(\eta,L)$, and the timescale in the gapless phase is much longer than that in the gapped phase. The system size is $L=18$ in both (a) and (b).
Memory loss effect.–
The spectral transition also leads to the transition of the timescale for the memory loss, i.e., $\tau_{\delta}(\eta,L)\simeq O(e^{L})$ in the gapless phase and $\tau_{\delta}(\eta,L)\simeq O(L^{0})$ in the gapped phase. Indeed, if we fix $V_{t}(\eta)$, |$\ket{\psi_{t}(\eta)}\simeq\ket{\Psi_{t,1}(\eta)}$ is independent of initial states |$\ket{\psi_{0}}$ for $t\geq\tau_{\delta}(\eta,L)$ within the precision $\delta$. Figure 3 shows $\bra{\psi_{t}(\eta)}X\ket{\psi_{t}(\eta)}$ with $X=\sum_{\ell}\sigma_{1}^{\ell}$ for several trajectories starting from different initial states, where $\sigma_{1}^{\ell}=\left(\bigotimes_{m=1}^{\ell-1}\sigma_{0}\right)\otimes\sigma_{1}\otimes\left(\bigotimes_{m=\ell+1}^{L}\sigma_{0}\right)$ is the $x$ component of the Pauli matrices at the site $\ell$. Here, $V_{t}(\eta)$ is generated through the Born rule based on one trajectory, and the same $V_{t}(\eta)$ is also applied to the other initial states. We can see that the expectation values corresponding to different initial states take almost the same value when $t\geq\tau_{\delta}(\eta,L)$ with $\delta=10^{-2}$, while different initial states lead to different expectation values when $t\ll\tau_{\delta}(\eta,L)$. We can understand that the timescale $\tau_{\delta}(\eta,L)$ of the memory loss effect for $\eta<\eta_{\mathrm{c}}$ is much larger than that for $\eta>\eta_{\mathrm{c}}$, due to the spectral transition.
We conjecture that the spectral gap, which becomes independent of trajectories in our monitored dynamics of pure states [44], takes the same value even when the initial state is a mixed state. This conjecture means that the spectral transition and thus the entanglement transition in the weakly monitored dynamics correspond to the purification transition: the volume-law (area-law) phase exhibits an exponentially long (constant) timescale in the system size for the initial mixed state to become a pure state approximately. This is consistent with the purification transition exposed to projective measurements [22], while the relation with the spectral transition has never been discussed before.
Structures of the entire spectrum.–
While we have focused on the spectral gap, we next compare the entire spectrum of the monitored dynamics with that of Hamiltonians in one-dimensional systems. Spectra of locally interacting many-body Hamiltonians usually have the following two structures: (a) the width of the spectrum becomes $O(L)$, which means that the typical level spacing is $O(e^{-L})$, and (b) the spectral gap becomes $O[1/\mathrm{poly}(L)]$ in gapless regimes or constant above a few states exponentially close to the ground state in gapped phases. We can analytically show that $\{\varepsilon_{i}(\eta)\}$ of our monitored dynamics also satisfy (a). Indeed, employing the majorization, the Horn theorem, and properties of doubly stochastic matrices [49], we can obtain [44]
$\varepsilon_{N}(\eta,L)-\varepsilon_{1}(\eta,L)\leq L\ln\left(\frac{1+\eta}{1-\eta}\right).$ (6)
On the other hand, the monitored dynamics do not satisfy (b) in the gapless phase; the Lyapunov spectra of low-lying excited states become $O(e^{-L})$, which is numerically illustrated in Supplemental Material [44].
Discussion.–
We have shown that the quantum dynamics exposed to weak measurements exhibit the spectral transition in terms of the Lyapunov spectrum, whose transition point coincides with that of the entanglement transition. This coincidence is analogous to ground-state phase transitions between gapless and gapped phases, while the scalings of the entanglement and the gap in the monitored dynamics are qualitatively different from those in isolated quantum systems. To establish the analogy and clarify the distinction between these transitions, exploring the locality of $K_{t}(\eta)$ should be an important future direction. While ground states in condensed matter models are usually determined from local Hamiltonians, it is non-trivial whether $K_{t}(\eta)$ becomes local. This distinction may be the origin of the difference in the scalings of the spectral gap and entanglement in gapless phases of these two transitions.
We have also shown that the spectral transition corresponds to the transition of the timescale for the memory loss of a pure state and conjectured that it also corresponds to the purification transition of mixed states. In monitored dynamics with projective measurements, where the purification transition is also reported [22], singular values of time-evolution operators can be strictly zero, and the Lyapunov spectrum with $i>1$ become ill-defined when the projective measurement is done to all qubits. In contrast, singular values $\{\Lambda_{t,i}(\eta)\}$ do not become zero in weak measurements, which may make relaxation dynamics under weak measurements distinct from those under projective measurements. Exploring the discrepancy and similarity between projective and weak measurements should be an interesting future work.
Acknowledgements.
We thank Yohei Fuji and Hisanori Oshima for fruitful discussions, Nobuyuki Yoshioka and Toshihiro Yada for valuable comments on the manuscript, and Zongping Gong for both. Derivation of the spectral gap under no unitary gate given in Supplemental Material is partly based on a note written by Toshihiro Yada. This work was supported by JST ERATO Grant Number JPMJER2302, Japan. K.M. was supported by JSPS KAKENHI Grant. No. JP23K13037. R.H. was supported by JSPS KAKENHI Grant No. JP24K16982.